ron has 63 books on 2 bookshelfs. the longer bookshelf has 5 more books than the shorter book shelf . how many book does the shorter bookshelf have?
Answer by ptaylor(2052)

Let x=number of books on shorter bookshelf
Then x+5=number of books on longer bookshelf
Now we are told that:
x+x+5=63 subtract 5 from both sides
x+x+5-5=63-5 collect like terms
2x=58 divide both sides by 2
x=29----------------------------number of books on shorter shelf-----ans
x+5=29+5=34--------------------------number of books on longer shelf
CK
29+34=63
63=63
also
29+5=34
34=34

Hope this helps----ptaylor